OpenStudy (anonymous):

28% of what number is 3 1/2 ?

OpenStudy (anonymous):

let the number be = x 0.28x = 3.5 x = 3.5/0.28 x = 12.5

OpenStudy (anonymous):

7/2 / 28/100 = 25/2 or 12 1/2